Apply Thevenins theorom to find o in the circuit below. VION 5AT) 161 50V
Answer #1

Problem Step! Calculation of Noh Convert Current Saree into 11 V aR = 16*5=804 ICR C + Apply kve in innerloor, - 80 of 16 ltster? Calculation of 4 RaH in shq (lety) are Connected in Paralle! som z Ron = 5x20 = 4m. 5+20 e Rah : Rih= lt4=5cm Thevinin